# LLM Honeypot: Leveraging Large Language Models as Advanced Interactive Honeypot Systems

### v1

@@ I Introduction
In parallel, recent advances in artificial intelligence and natural language processing have given rise to Large Language Models (LLMs) capable of generating human-like text responses [8]. With appropriate fine-tuning and prompt engineering, these models have the potential to revolutionize honeypot technology by enabling the creation of highly realistic and interactive honeypots. Leveraging LLMs, honeypots can engage with attackers in a more sophisticated manner, providing more secure and intelligent responses. Recent research demonstrates the feasibility of using LLMs as dynamic honeypot servers by employing pre-trained models like ChatGPT. Even without extensive fine-tuning, well-crafted prompts can allow these models to observe and study attacker behaviors and tactics effectively [9, 10, 11]. However, a significant challenge in using chatbots as honeypots is the potential for attackers to detect and identify the honeypot due to static elements or predictable behaviors. One way to fix this problem would be to make honeypot environments more dynamic and real, use advanced behavioral analysis, and create continuous learning models that can better adapt to new attack patterns [9]. These advancements would enhance the capability of honeypots to remain covert and effective in detecting sophisticated threats.
Therefore, in this study, we aim to develop an interactive LLM-Honeypot system based on Large Language Models (LLMs) capable of accurately mimicking the behavior of a Linux server. By fine-tuning a pre-trained open-source language model using a dataset of attacker-generated commands and responses, we seek to create an interactive honeypot system that can engage with attackers in a highly realistic and informative manner. This approach has the potential to significantly enhance the effectiveness of honeypot technology, offering cybersecurity professionals a powerful tool to detect and analyze malicious activities.
@@ II Methodology
The fine-tuned model is then rigorously evaluated to ensure it effectively engages with attackers and provides valuable security insights. Finally, the optimized honeypot is deployed to a public IP address for real-world interaction with potential threats.
@@ II-B Prompt Engineering
For the second part of the dataset, we created a set of prompts that positioned the model as a Linux expert, capable of providing detailed explanations of Linux terminal commands. These prompt enabled the model to generate informative and context-rich responses that demonstrated a deep understanding of Linux commands and their applications. Again, a sample prompt is provided in Listing 2.
@@ II-C Model Selection
We tested several recent models, including Llama3 [17], Phi 3 [18], CodeLlama [19], and Codestral [20]. Llama3, with its 8B and 70B variants, offers scalable language processing, while Phi 3, CodeLlama, and Codestral are notable for their focus on code-development tasks. However, our experiments showed that code-centric models like Phi 3, CodeLlama, and Codestral were less effective for honeypot simulation. Larger models (8B and 70B) were too slow, emphasizing the need for computational speed. Smaller models demonstrated sufficient capability, suggesting the importance of balancing model size and efficiency.
@@ III-B Custom SSH Server Wrapper
To deploy the final model as a functional honeypot server, we crafted a custom SSH server using Python’s Paramiko library [25]. This server integrates our fine-tuned language model to generate realistic responses.
@@ III-D Similarity Analysis with Cowrie Outputs
Figure 5: Histogram of Cosine Similarity Scores over 140 Samples
@@
It is also worth noting that while some generated outputs may differ but they remain contextually accurate and true for the given commands.
@@ IV Conclusion
Future work includes expanding training datasets, exploring alternative fine-tuning, and incorporating behavioral analysis. We plan to deploy the system to a public IP for real-world interaction, collecting logs on attack vectors for analysis. This will involve creating knowledge graphs to map attack techniques and provide insights into attacker strategies. Performance analysis can focus on response realism, engagement effectiveness, and robustness using metrics like accuracy and interaction quality, aiming to refine our model and enhance honeypots for better cyber-threat detection and analysis.

### v2

@@ I Introduction
In parallel, recent advances in artificial intelligence and natural language processing have given rise to Large Language Models (LLMs) capable of generating human-like text responses [8]. With appropriate fine-tuning and prompt engineering, these models have the potential to revolutionize honeypot technology by enabling the creation of highly realistic and interactive honeypots. Leveraging LLMs, honeypots can engage with attackers in a more sophisticated manner, providing more secure and intelligent responses. Recent research demonstrates the feasibility of using LLMs as dynamic honeypot servers by employing pre-trained models like ChatGPT. Even without extensive fine-tuning, well-crafted prompts can allow these models to observe and study attacker behaviors and tactics effectively [9, 10, 11]. However, a significant challenge in using chatbots as honeypots is the potential for attackers to detect and identify the honeypot due to static elements or predictable behaviors. One way to fix this problem would be to make honeypot environments more dynamic and real, use advanced behavioral analysis, and create continuous learning models that can better adapt to new attack patterns [9].
Our approach differs from previous LLM Honeypot studies [9, 10, 11] by utilizing supervised fine-tuning, while earlier research relied on prompt engineering and storing conversation histories as context. For example, [11] used chain-of-thought and few-shot prompting to enhance LLM honeypots’ realism, while [9] appended all past commands and outputs for model completion, though limited by ChatGPT’s 8,000-token cap. Unlike previous work using closed-source LLMs like ChatGPT, our research develops an interactive honeypot system using a fine-tuned, open-source LLM. By training it on attacker-generated commands, we aim to create a system that mimics Linux server behavior and engages attackers more realistically, enhancing honeypot effectiveness for cybersecurity professionals.
@@ II Methodology
The fine-tuned model is rigorously evaluated to ensure it effectively engages with attackers and provides valuable security insights. Finally, the optimized honeypot is deployed to a public IP address for real-world interaction with potential threats.
@@ II-B Prompt Engineering
For the second part of the dataset, we created prompts that positioned the model as a Linux expert, capable of offering detailed explanations of terminal commands. These prompts helped the model gain a deeper understanding of Linux commands and their applications. A sample prompt is provided in Listing 2. The dataset was used to improve the model’s ability to summarize commands, aiming for more accurate and contextually relevant responses.
@@ II-C Model Selection
We tested several recent models, including Llama3 [17], Phi 3 [18], CodeLlama [19], and Codestral [20]. Llama3, with its 8B and 70B variants, offers scalable language processing, while Phi 3, CodeLlama, and Codestral are notable for their focus on code-development tasks. However, our experiments showed that code-centric models were less effective for honeypot simulation. Larger models (70B) were too slow, emphasizing the need for computational speed. Smaller models (8B) demonstrated sufficient capability, suggesting the importance of balancing model size and efficiency.
@@ III-B Custom SSH Server Wrapper
To deploy the final model as a functional honeypot server, we crafted a custom SSH server using Paramiko library [25]. This server integrates our fine-tuned language model to generate realistic responses.
@@ III-D Similarity Analysis with Cowrie Outputs
Some outputs may vary but remain contextually accurate and true to the commands. The LLM honeypot server consistently handles out-of-bound or invalid commands due to strict system prompt guardrails and training on erroneous examples. For instance, when encountering unrecognized commands, the model replicates realistic system behavior with messages like ‘bash: XXX: command not found.’ Testing shows the model maintains its persona and provides convincing responses, even with unexpected inputs.
Figure 5: Histogram of Cosine Similarity Scores over 140 Samples